OpenStudy (anonymous):

if csc3a=sec9a, the m

OpenStudy (anonymous):

i think i stumped ya'll hahaha

OpenStudy (dumbcow):

you can rewrite this as: \[\sin(3a) = \cos(9a)\] then you can relate cos to sin by shifting cos to right by pi/2 \[\rightarrow \cos(9a) = \sin(9a+\pi/2)\] thus \[\sin(3a) = \sin(9a + \pi/2)\] \[3a = 9a + \pi/2\] a = -pi/12

OpenStudy (anonymous):

lol you just pulled that outta your butt didnt you?

OpenStudy (anonymous):

it says the answer is 7.5, its just I have no idea how to get it

OpenStudy (dumbcow):

kinda..but it works

OpenStudy (anonymous):

lol it didnt get the answer though any ideas?

OpenStudy (anonymous):

this has got to be the dumbest question ever

OpenStudy (dumbcow):

there are multiple answers to this question

OpenStudy (dumbcow):

you can also say \[\cos(9a) = \sin(\pi/2 - 9a)\] then 3a = pi/2 - 9a a = pi/24 = 7.5 degrees

OpenStudy (anonymous):

oh alright dang that confusing

myininaya (myininaya):

 \[\sin(3a)=\cos(9a)\] \[\cos(9a)=\sin(\frac{\pi}{2}-9a)\] So we have \[\sin(\frac{\pi}{2}-9a)=\sin(3a)\] But we also have \[\sin(3a)=\sin(3a+2n \pi) \] Though we could say the same about the other side But anyways we have \[\sin(\frac{\pi}{2}-9a)=\sin(3a+2 n \pi)\] \[\frac{\pi}{2}-9a=3a+2 n \pi\] solve for a